# Identifying and Addressing Delusions for Target-Directed Decision Making

### v4

@@ Abstract
Target-directed agents utilize self-generated targets, to guide their behaviors for better generalization. These agents are prone to blindly chasing problematic targets, resulting in worse generalization and safety catastrophes. We show that these behaviors can be results of delusions, stemming from improper designs around training: the agent may naturally come to hold false beliefs about certain targets. We identify different types of delusions via intuitive examples in controlled environments, and investigate their causes and mitigations. With the insights, we demonstrate how we can make agents address delusions preemptively and autonomously. We validate empirically the effectiveness of the proposed strategies in correcting delusional behaviors and improving out-of-distribution generalization.
@@ 1 Introduction
Thus, a certain type of decision-time planning agents are developed, which make use of their more adaptive planning outcomes, which we call targets, to direct its behaviors to generalize better in novel situations (Alver & Precup, 2022). For instance, these targets can help decompose a complicated given task into smaller and more familiar steps. In this paper, we use targets to denote (sub-)goals produced by the agents themselves during decision-time planning. These “target-directed” agents make use of generators to propose candidate targets, as well as optional estimators to evaluate the favorability of the candidates, to assist the selection of a target to commit to.
While target-directed agents were supposed to be able to generalize well OOD, they are often observed to be blindly chasing unreachable or unsafe targets (Bengio et al., 2024). While existing literature has largely neglected the investigations into these behaviors, we introduce the readers to a new perspective to show that these agents are likely (unintentionally) designed to be delusional. Unlike hallucinations, delusions are obviously wrong beliefs that an agent holds, and reflect an inability to reject false beliefs (Kiran & Chaudhury, 2009). These in the RL context correspond to false beliefs that are natural results of an agent’s improper learning process, which limit the agents’ situational understanding and lead to worse generalization or even safety catastrophes. Delusional behaviors stem from the incoordination of the generator and the estimator: problematic targets should not be generated, and even if they were, they should not be favored (Corlett, 2019).
This paper focuses on necessary designs for an agent to address its delusions autonomously and preemptively: 1) incorporation of an estimator, which enables the rejection of problematic targets; 2) adequate learning rules with which false beliefs about targets can be counteracted and 3) proper training data that include those counteracting delusions. With clear diagnoses in the designed environments, we identify types of delusions, and investigate more proper designs of the training processes. We then applied our idea to hindsight relabeling, a popular and simple form of target-directed RL training and validated in experiments that our strategies led to lower delusion-related errors and significant reduction in delusional behaviors, resulting in significant improvements in OOD generalization performance.
@@ 2 Preliminaries
Source-Target Pairs & Hindsight Relabeling Essentially, target-directed agents are trained on “source-target pairs”, which are organized training data that could let the agents learn about the relationship between one state and another. Training target-directed agents with only contemporary targets (the ones being followed) can lead to poor performance (Dai et al., 2021), since contemporary targets may be low-quality, hard to achieve, or lack in diversity (Moro et al., 2022; Davchev et al., 2021). Hindsight Experience Replay (HER) is proposed to be an experience organization mechanism transforming collected experiences into relabeled transitions augmented with alternative targets (Andrychowicz et al., 2017). HER augments a transition $\langle x_{t},a_{t},r_{t+1},x_{t+1}\rangle$ with an additional observation $x^{\odot}$ (or its encoding), the relabeled target, creating a source-target pair, with one decision point for the current step and another for the relabeled target. Relabeling strategies, which correspond to how $x^{\odot}$ is selected, are critical for the performance of HER-trained agents (Shams & Fevens, 2022). Most popular choices are trajectory-level, meaning $x^{\odot}$ comes from the same trajectory as $x_{t}$. These include “future”, where $x^{\odot}=x_{t^{\prime}}$ with $t^{\prime}>t$, and “episode”, with $0\leq t^{\prime}\leq T_{\perp}$.
@@
SSM In existing works, failure modes of target-directed agents are often overlooked, possibly due to a lack of access to environmental ground truth in benchmark environments. To identify the causes, and provide intuitive examples, we craft a set of fully-observable environments based on the MiniGrid-BabyAI framework (Chevalier-Boisvert et al., 2018b; Hui et al., 2020), named SwordShieldMonster (SSM for short). In SSM, the agent moves one step at a time in $4$ absolute directions to navigate fields with episode-terminating, randomly placed lava traps. The density of traps is specified by a difficulty parameter $\delta$, while guaranteeing a viable path to success, i.e. acquiring a sparse terminal reward. Agents must collect a sword and a shield, both also randomly placed, before reaching the “monster” guarding the treasure. Reaching the monster prematurely ends the episode without reward. Visualizations of SSM are in Fig. 2, with examples of delusional behaviors discussed later. The sword and shield are picked up by moving to their respective grid cells, and agents cannot drop them. Such design introduces temporary unreachability in the state structure, as non-terminal states are not fully traversable from one to another. Semantically, this segments SSM states into $4$ equivalence classes, formed by $2$ binary indicators: the agent’s possession of the sword and shield. For example, $\langle 0,1\rangle$ denotes sword not acquired, shield acquired”. Despite SSM’s simplicity in terms of observation space, a similar set of environments much simpler than SSM was already shown to be a significant challenge for state-of-the-art methods due to the OOD challenges (Zhao et al., 2024).
@@ 3 Delusions in Target-Directed RL
To ultimately make the agents able to address delusions autonomously, we first need to identify where the problems reside. We first look into the generator to find its contributions to the delusional behaviors, and then locate delusions and their causes in the estimator.
(a) E.3.2.1 (current state in $\langle 1,1\rangle$)
(b) E.3.2.2 (current state in $\langle 1,1\rangle$, target in $\langle 0,0\rangle$)
Figure 2: Visualizations of Delusional Behaviors in SSM: The agent location is marked as a red triangle, lava traps’ as orange squares, and monsters’ as green loops. In both a) & b), the estimators, lacking exposure to problematic targets during training, failed to estimate their reachability. This led to the delusion that these problematic targets (whose location are marked with yellow dots), seemingly close to every state, can be used as proxies to reach the distant task goal almost instantly.
3.1 Generator’s Role in Delusional Behaviors: Bad Targets
Most learned generators are subjected to unwanted generalization, thus will inevitably generate problematic targets, a behavior sometimes characterized as hallucination (Jafferjee et al., 2020). While hallucination can be reduced to a degree, they are generally unavoidable, especially in OOD scenarios. Despite that not all delusions are caused by problematic targets, generator hallucinations pose a major risk to safety.
@@ 3.1 Generator’s Role in Delusional Behaviors: Bad Targets
The second type includes targets that correspond to valid states in the task MDP, but cannot be fulfilled from the current state. Unlike G.3.1, these targets are only temporarily inappropriate and could be perfectly good targets if the agent were currently to be in a different state. There are some notable subtypes, such as: 1) Irreversible: targets once reachable but are now blocked due to a past transition. For instance, after acquiring the sword in SSM, the agent transitions from class $\langle 0,0\rangle$ to $\langle 1,0\rangle$, sealing off access to $\langle 0,0\rangle$ or $\langle 0,1\rangle$; 2) Segregated: targets that can only be reached from specific initial states that are not available to the agent. For example, an agent spawned in class $\langle 1,0\rangle$ cannot reach states in $\langle 0,0\rangle$ or $\langle 0,1\rangle$.
G.3.1.1 targets are often overlooked in literature, since hallucinations are mostly discussed in contexts without temporal progression, and also they do not exist in all MDPs. G.3.1.1 can appear more frequently if we do not intentionally train the generators to avoid generating them. For example, a conditional target generator which learns from “episode” will be more likely to produce G.3.1.1 targets, such as the one in Fig. 2 b). This applies to training beyond HER as well.
@@ 3.2 Estimator’s Role in Delusional Behaviors: Bad Evaluation
Incorporating an estimator does not mean delusions addressed, as the estimators’ own delusional beliefs about targets can directly cause delusional behaviors. A lack of either of the following two aspects can result in blind spots of knowledge about targets, i.e. delusions regarding both delusional or non-delusional targets.
@@
Training Data Even with capable update rules, delusions can still arise from the lack of exposure to data counteracting the estimators’ false beliefs. This can often be caused by improper experience organization mechanisms, which are responsible for transforming the experienced interactions with environments into training data, as shown in Fig. 1. Despite these mechanisms like HER can lead to more diversity in training data, they are often not capable of avoiding delusions: 1) Certain relabeling strategies naturally cause delusions. For instance, “future” only relabels with future observations, thus only exposes a learner to future reachable targets, leaving the estimator to guess when a “past” target is proposed at decision-time; 2) Trajectory-level relabeling can also be problematic. Short trajectories, common in many training procedures, only cover limited portions of the state space and prevent estimators from learning about distant targets, leaving estimators to guess when a distant target is generated at decision-time. Short trajectories can be a product of experimental design (initial state distributions, Maximum Episode Lengths, MELs (Erraqabi et al., 2021)) or innate environmental characteristics (density of terminal states). It is worth noting that delusions caused by improper training data (organizing mechanisms) also extend to cases like few-shot generalization.
@@
The standalone type of delusions includes false estimations about non-delusional targets. E.3.2 delusions can lead to undesirable or delusional targets being favored. E.3.2(-induced) behaviors can hurt the agents’ generalization abilities.
@@
4 Addressing Delusions: Strategies Applied to Hindsight Relabeling
Identifying delusions enables mitigation strategies. We develop methodologies applicable to general target-directed agents coming from a spectrum of training procedures. Then, we provide examples on how we can materialize such methodologies in agents trained with HER, because of its popularity and the generality behind hindsight relabeling (and source-target pairs).
Assume we are dealing with a hypothetical target-directed agent with a dual-component architecture as in Fig. 1, learning both components exclusively from hindsight-relabeled transitions. The generator proposes targets corresponding to potentially distant states, and the estimator learns the favorability of targets by (continuously) punishing the unachieved or unsafe targets present in the training data. As discussed earlier, to address delusions, we need effective update rules and training data distribution. By assuming this framework, we skip discussing update rules, as they depend on the agents’ specific designs. Our emphasis on the training data reflects the fact that most existing RL agents only learn from experienced data, while addressing delusions requires learning from targets that can never be experienced. This corresponds to a discrepancy between target-directed agents’ behaviors and training: at decision-time, targets outside experience can be proposed; while during training, only experienced targets are learned from.
@@ 4.1 Assistive Strategies for Addressing Delusions
We first introduce two strategies to improve training data distributions, which are then materialized as two hindsight relabeling strategies for HER. Later, we compare the ideas behind the proposed strategies with existing ones, and study how the two groups can be combined to create mixtures for estimators and hybrids for the generator-estimator duo.
The following strategies seek to expand the support of the training data distribution, to include those source-target combinations that the agent could never experience, i.e. those involving G.3.1 and G.3.1.1 targets. With the assumed update rules, this extra exposure is expected to address E.3.2.1 and E.3.2.2 estimator delusions.
@@
In Zhao et al. (2024), the authors identified a delusional behavior resulting from E.3.2.1 delusions, in the language of this paper, and they proposed to train the estimator additionally with candidate targets proposed by the generator. With HER, we can transform this auxiliary loss into a Just-In-Time (JIT) HER strategy that, when a transition is sampled for training, relabels it with a conditionally generated target by the generator. We can expect “generate” to be effective for training estimators, as the estimators will get exposed to all kinds of problematic targets that the generator could offer: When a generator learns to generate mostly viable targets, “generate” will be primarily effective at dealing with the E.3.2 and E.3.2.1 (G.3.1s are unlikely to be eradicated). “generate” requires the use of the generator, thus it incurs additional computational burden, depending on the complexity of target generation processes.
Adapting “generate” beyond HER should be straightforward: whenever a target is involved during estimator learning, we may replace it with generated targets.
@@
The second strategy, namely “pertask”, is to expose the estimator to targets that were experienced before, s.t. it could figure out that some previously achieved targets are not achievable from the current state.
We materialize “pertask” into an assistive atomic relabeling strategy that relabels transitions with observations from the same training task, sampled across the entire memory. “pertask” works on a task level, creating more diverse source-target pairs than trajectory-level strategies like “future” & “episode”, which bias the training data distribution toward shorter-distance source-target pairs. This approach additionally brings exposure to the estimator to learn against E.3.2.2 delusions and to the long-distance source-target pairs against E.3.2 caused by short trajectories.
It is worth noting that “pertask” also biases the training data distribution, making the agent spread out its efforts into learning the source-target pairs that are potentially far away from each other. Despite the increment to training data diversity, long-distance pairs are less likely to contribute to better decision-making compared to the shorter-distance in-episode ones offered by “episode”, as shown later in experiments.

### v5

# Identifying and Addressing Delusions for Target-Directed Decision-Making

@@ Abstract
Target-directed agents utilize self-generated targets, to guide their behaviors for better generalization. These agents are prone to blindly chasing problematic targets, resulting in worse generalization and safety catastrophes. We show that these behaviors can be results of delusions, stemming from improper designs around training: the agent may naturally come to hold false beliefs about certain targets. We identify delusions via intuitive examples in controlled environments, and investigate their causes and mitigations. With the insights, we demonstrate how we can make agents address delusions preemptively and autonomously. We validate empirically the effectiveness of the proposed strategies in correcting delusional behaviors and improving out-of-distribution generalization.
@@ 1 Introduction
Thus, embracing the concept of intentions, a certain type of decision-time planning agents are developed, which make use of their more adaptive planning outcomes, which we call targets, to direct its behaviors to generalize better in novel situations (Alver & Precup, 2022). For instance, these targets can help decompose a complicated task into small and familiar steps. In this paper, we use targets to denote (sub-)goals produced by the agents themselves during decision-time planning. These “target-directed” agents make use of generators to propose candidate targets, as well as optional estimators to evaluate the favorability of the candidates, to select a target to commit to.
While target-directed agents were supposed to be able to generalize well OOD, they are often observed to be blindly chasing unreachable or unsafe targets (Bengio et al., 2024). This paper focuses on the investigations of these behaviors, which existing literature has largely neglected. We introduce the readers to a new perspective to show that these agents are likely designed to be delusional. Unlike hallucinations, delusions are obviously wrong beliefs that an agent holds, and reflect an inability to reject false beliefs (Kiran & Chaudhury, 2009). These in the RL context correspond to false beliefs that are natural results of an agent’s improper learning process, which limit the agents’ situational understanding and lead to worse generalization or even safety catastrophes. Delusional behaviors stem from the incoordination of the generator and the estimator: problematic targets should not be generated, and even if they were, they should not be favored (Corlett, 2019).
This paper discusses some necessary conditions for agents to address delusions autonomously and preemptively: 1) incorporation of an estimator, enabling the rejection of problematic targets; 2) adequate learning rules with which false beliefs about targets can be counteracted and 3) proper training data that can counteract delusions. With clear diagnoses in the controlled environments, we identify types of delusions, and investigate proper designs of the training processes. We then applied our ideas to hindsight relabeling, a fundamental form of target-directed RL training, and validated experimentally that our strategies led to lower delusion-related errors and reduction in delusional behaviors, resulting in significant improvements in OOD generalization performance.
@@ 2 Preliminaries
Source-Target Pairs & Hindsight Relabeling Essentially, target-directed agents are trained on “source-target pairs”, which are organized training data that could let the agents learn about the relationship between the source state and the target state. Training target-directed agents with only contemporary targets (the ones being followed) can lead to poor performance (Dai et al., 2021), since contemporary targets may be low-quality, hard to achieve, or lack in diversity (Moro et al., 2022; Davchev et al., 2021). Hindsight Experience Replay (HER) is proposed to be an experience organization mechanism, transforming collected experiences into relabeled transitions augmented with alternative targets (Andrychowicz et al., 2017). HER augments a transition $\langle x_{t},a_{t},r_{t+1},x_{t+1}\rangle$ with an additional observation $x^{\odot}$ (or its encoding), the relabeled target, creating a source-target pair, with one decision point for the current step and another for the relabeled target. Relabeling strategies, which correspond to how $x^{\odot}$ is selected, are critical for the performance of HER-trained agents (Shams & Fevens, 2022). Most popular choices are trajectory-level, meaning $x^{\odot}$ comes from the same trajectory as $x_{t}$. These include “future”, where $x^{\odot}=x_{t^{\prime}}$ with $t^{\prime}>t$, and “episode”, with $0\leq t^{\prime}\leq T_{\perp}$.
@@
SSM In existing works, failure modes of target-directed agents are often overlooked, possibly due to a lack of access to ground truths in benchmark environments. To identify the causes, and provide intuitive examples, we craft a set of fully-observable environments based on the MiniGrid-BabyAI framework (Chevalier-Boisvert et al., 2018b; Hui et al., 2020), named SwordShieldMonster (SSM for short). In SSM, the agent moves one step at a time in $4$ absolute directions to navigate fields with episode-terminating, randomly placed lava traps. The density of traps is specified by a difficulty parameter $\delta$, while guaranteeing a viable path to success, i.e. acquiring a sparse terminal reward. Agents must collect a sword and a shield, both also randomly placed, before reaching the “monster” guarding the reward. Reaching the monster prematurely ends the episode. Visualizations of SSM are in Fig. 2, with examples of delusional behaviors discussed later. The sword and shield are picked up by moving to their respective grid cells, and agents cannot drop them. Such design introduces temporary unreachability in the state structure, as non-terminal states are not fully traversable from one to another. Semantically, this segments SSM states into $4$ equivalence classes, formed by $2$ binary indicators: the agent’s possession of the sword and shield. For example, $\langle 0,1\rangle$ denotes sword not acquired, shield acquired”. Despite SSM’s simplicity in terms of observation space, a similar set of environments much simpler than SSM was already shown to be a significant challenge for state-of-the-art methods due to the OOD challenges (Zhao et al., 2024).
@@ 3 Delusions in Target-Directed RL
To ultimately enable the agents to address delusions autonomously, we first need to identify where the problems reside. We first look into the generator to find its contributions to the emergence of delusional behaviors, and then locate delusions and their causes in the estimator.
(a) E.3.2.1: both planned steps are delusional
(b) E.3.2.2 (current state in $\langle 1,1\rangle$, target in $\langle 0,0\rangle$): first planned step is delusional
Figure 2: Delusional Behaviors in SSM: The agent location is marked as a red triangle, lava traps’ as orange squares, and monsters’ as green loops. In both cases, the estimators, lacking understanding of the problematic targets, fail to estimate the reachability of the problematic targets proposed at decision-time (yellow dots), leading to the delusions that shorter paths to the task goal exist through the problematic targets.
3.1 Generator’s Role in Delusional Behaviors: Proposing Bad Targets
Most learned generators are subjected to unwanted generalization, thus inevitably generate problematic targets (unreachable or unsafe), a behavior characterized as hallucination (Jafferjee et al., 2020). While hallucination can be reduced to a degree, they are generally unavoidable, especially in OOD scenarios. Despite that not all delusions are caused by problematic targets, generator hallucinations pose a major safety risk (Liu et al., 2022; Bengio et al., 2024).
@@ 3.1 Generator’s Role in Delusional Behaviors: Proposing Bad Targets
The second type includes targets that correspond to valid states in the task MDP, but cannot be fulfilled from the current state. Unlike G.3.1, these G.3.1.1 targets are only temporarily inappropriate and could be perfectly good targets if the agent were in a different state. For the time-dependency, G.3.1.1 can be easily overlooked. There are some notable subtypes, such as: 1) Irreversible: targets once reachable but are now blocked due to a past transition. For instance, after acquiring the sword in SSM, the agent transitions from class $\langle 0,0\rangle$ to $\langle 1,0\rangle$, sealing off access to $\langle 0,0\rangle$ or $\langle 0,1\rangle$; 2) Segregated: targets that can only be reached from specific initial states that are not available to the agent. For example, an agent spawned in class $\langle 1,0\rangle$ cannot reach states in $\langle 0,0\rangle$ or $\langle 0,1\rangle$.
G.3.1.1 targets are often overlooked in literature, since hallucinations are mostly discussed in contexts without temporal progression, and also they do not exist in all MDPs. G.3.1.1 can appear more frequently if we do not avoid training the generators with them. For example, a conditional target generator which learns from “episode” will be more likely to produce G.3.1.1 targets (compared to “future”), such as the one in Fig. 2 b). This applies to training beyond HER as well.
@@ 3.2 Estimator’s Role in Delusional Behaviors: Bad Evaluation
Incorporating an estimator only delegates the responsibility of avoiding delusional behavior from the generator to the estimator, as the estimators’ own delusional beliefs about targets can directly cause delusional behaviors. A lack of either of the following two necessary aspects results in blind spots of knowledge about targets, i.e. delusions regarding both delusional or non-delusional targets.
@@
Training Data Even with proper update rules, delusions can still arise from the lack of exposure to data counteracting the estimators’ false beliefs. This is often caused by improper experience organization mechanisms, responsible for transforming the experienced interactions with environments into training data, as shown in Fig. 1. Despite these mechanisms like HER can lead to more diversity in training data, they are often accompanied by delusions: 1) Certain relabeling strategies naturally cause delusions. For instance, “future” only relabels with future observations, thus only exposes a learner to future reachable targets, leaving the estimator to guess when a “past” target is proposed at decision-time; 2) Trajectory-level relabeling can also be problematic. Short trajectories, common in many training procedures, only cover limited portions of the state space and prevent estimators from learning about distant targets, leaving estimators to guess when a distant target is generated at decision-time. Short trajectories can be a product of experimental design (initial state distributions, Maximum Episode Lengths, MELs (Erraqabi et al., 2021)) or innate environmental characteristics (density of terminal states). It is worth noting that delusions caused by improper training data (organizing mechanisms) also extend to cases like few-shot generalization.
@@
The standalone type of delusions describes false estimations about non-delusional targets. E.3.2 delusions can lead to undesirable or delusional targets being favored and thus can hurt generalization.
@@
4 Addressing Delusions in Target-Directed Agents
Having identified delusions, taking advantage of the shared reliance on source-target pairs, we can develop strategies applicable to general target-directed agents coming from various training procedures. Then, we provide examples on how to materialize such strategies in agents trained with HER, because of its popularity, its direct association with source-target pairs and the fact that one can trivially transform HER-based training into losses established over two sampled transitions from ordinary experience replays.
Without the loss of generality, assume we are dealing with a hypothetical target-directed agent with a dual-component architecture as in Fig. 1, learning both components exclusively from hindsight-relabeled transitions. The generator proposes targets corresponding to potentially distant states, and the estimator learns the favorability of targets by (continuously) punishing the unachieved or unsafe targets sampled during training. Note that even for methods whose generators are not trained with HER, our estimator-focused strategies still apply.
As discussed earlier, to address delusions, we need effective update rules and training data. By assuming this framework, we skip discussing update rules, as they depend on the specific designs of the chosen target-directed agents. Our emphasis on the training data shines light on the fact that most existing RL agents only learn from experienced data, while addressing delusions requires learning from targets that can never be experienced. This corresponds to a discrepancy between (most existing) target-directed agents’ behaviors and training: at decision-time, targets outside experience can be proposed; while during training, only experienced targets are learned from.
@@ 4.1 Assistive Strategies for Addressing Delusions
We first introduce two ideas to improve training data distributions, which can then be materialized as two hindsight relabeling strategies for HER. These ideas seek to expand the support of the training data distribution, to include those source-target combinations that the agent could never experience, i.e. those involving G.3.1 and G.3.1.1 targets. Since our proposed ideas do not rely on additional assumptions, they should be expected to be applicable generally. Additionally, with convergent learning rules (provided by the target-directed framework that these strategies are applied to), the strategies should also lead to the correct estimation of G.3.1 and G.3.1.1 targets.
Later, we compare the ideas behind all HER strategies, including the existing and the newly proposed, and study how the two groups can be combined to accommodate the learning needs.
@@
Zhao et al. (2024) identified delusional behaviors resulted from E.3.2.1 delusions, in the language of this paper, and proposed to train the estimator additionally with candidate targets proposed by the generator. With HER, we can transform this auxiliary loss into a Just-In-Time (JIT) HER strategy that, as a transition is sampled for training, relabels it with a proposed target by the generator. We can expect “generate” to be effective for training estimators, as the estimators will get exposed to all kinds of problematic targets that the generator could offer: When a generator learns to generate mostly viable targets, “generate” will be primarily effective at dealing with the E.3.2 and E.3.2.1 (G.3.1s are unlikely to be eradicated). “generate” requires the use of the generator, thus it incurs additional computational burden, depending on the complexity of target generation processes.
Adapting “generate” beyond HER should be straightforward: whenever a target is involved during estimator learning, we may replace it with generated targets. “generate” comes at a cost of extra computation, to which attention may be required in real-world applications with speed demands.
@@
The second strategy, namely “pertask”, is to expose the estimator to all targets experienced before, s.t. it could figure out that some previously achieved targets are unreachable from the current state.
We materialize “pertask” into an assistive atomic strategy that relabels transitions with observations from the same training task, sampled across the entire memory. Importantly, “pertask” brings exposure to the estimator to learn against E.3.2.2 delusions and to the long-distance source-target pairs against E.3.2 caused by short trajectories. Take SSM as an intuitive example, a current state in situation $\langle 1,0\rangle$ in the current episode can now be paired with a target in $\langle 0,1\rangle$ in another episode, s.t. the estimator learns such G.3.1.1 target is unreachable, as shown in Fig. 4 (in Appendix).
It is worth noting that “pertask” also biases the training data distribution, making the agent spread out its efforts into learning the source-target pairs potentially far away from each other. Despite increasing training data diversity, long-distance pairs are less likely to contribute to better decision-making compared to the shorter-distance in-episode ones offered by “episode”, as shown later.